OpenStudy (anonymous):

What is the kinetic energy of an automobile with a mass of 1210 kg traveling at a speed of 13.0 m/s?

OpenStudy (anonymous):

Ek = 0.5*m*v^2 use this equation hope so it will help u out

OpenStudy (anonymous):

KE = 1/2mv^2 So therefore KE = 0.5(1210kg)(13 m/s)^2